jason78687:

A system of equations consisting of a linear equation and a quadratic equation has infinitely many solutions. Is this statement true? Always Sometimes Never

NotLeiland:

I'm not sure how to explain it to you but yes, it is always true.

jason78687:

@notleiland wrote:
I'm not sure how to explain it to you but yes, it is always true.
Actually i was stuck between sometimes and never no way always is true

NotLeiland:

@jason78687 wrote:
@notleiland wrote:
I'm not sure how to explain it to you but yes, it is always true.
Actually i was stuck between sometimes and never no way always is true
Expert-Verified Answer Given : A system of equations consisting of a linear equation and a quadratic equation has infinitely many solutions .Jul 1, 2020 https://brainly.in/question/19021967#:~:text=Expert%2DVerified%20Answer&text=Given%20%3A%20A%20system%20of%20equations,equation%20has%20infinitely%20many%20solutions.

thatgirlcrazy:

the statement could never be true because As a linear Equation can cut Quadratic equation maximum at two points hence maximum possible solutions are two so its not possible to have infinite solution Quadratic Equation : f(x) = ax² + bx + c Linear equation f(x) = mx + d Equating both ax² + bx + c = mx + d => ax² + x(b-m) + c - d = 0 a form of Quadratic equation hence maximum possible solutions are 2 Hence A system of equations consisting of a linear equation and a quadratic equation has infinitely many solutions is Never True

jason78687:

@notleiland wrote:
@jason78687 wrote:
@notleiland wrote:
I'm not sure how to explain it to you but yes, it is always true.
Actually i was stuck between sometimes and never no way always is true
Expert-Verified Answer Given : A system of equations consisting of a linear equation and a quadratic equation has infinitely many solutions .Jul 1, 2020 https://brainly.in/question/19021967#:~:text=Expert%2DVerified%20Answer&text=Given%20%3A%20A%20system%20of%20equations,equation%20has%20infinitely%20many%20solutions.
I dont trust this they are wrong a lot of times

NotLeiland:

@jason78687 wrote:
@notleiland wrote:
@jason78687 wrote:
@notleiland wrote:
I'm not sure how to explain it to you but yes, it is always true.
Actually i was stuck between sometimes and never no way always is true
Expert-Verified Answer Given : A system of equations consisting of a linear equation and a quadratic equation has infinitely many solutions .Jul 1, 2020 https://brainly.in/question/19021967#:~:text=Expert%2DVerified%20Answer&text=Given%20%3A%20A%20system%20of%20equations,equation%20has%20infinitely%20many%20solutions.
I dont trust this they are wrong a lot of times
You do you I guess